OpenStudy (anonymous):

please help 4x + 5 = 29

OpenStudy (anonymous):

4x+5=29 -5 -5 4x=24 x=6

OpenStudy (mathstudent55):

Two step equation: first, subtract 5 from both sides. Then, divide both sides by 4.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

welcome!

OpenStudy (skullpatrol):

The side containing the variable involves a certain order of operations apply the inverse operations in the opposite order.